The particular integral of (x2D2-xD+4)Y=COS(LOGX)is
1
Expert's answer
2021-08-23T16:22:26-0400

"(x^2D^2-xD+4)y=cos(logx)\\
\text{By euler equation.}\\
(D'(D'-1)-D'+4)y=cosz\\
(D'^2-2D'+4)y=cosz\\
\text{homogeneous differential equation is given by}\\
(D'^2-2D'+4)y=0\\
\text{Auxiliary equation,}\\
m^2-m+4=0\\
m=1\pm\sqrt3i\\
\text{Therefore, general solution of homogeneous differential equationis}\\y=e^z(c_1cos(\sqrt3z)+c_2sin(\sqrt3z))\\
y=x(c_1cos(\sqrt3logx)+c_2sin(\sqrt3logx))\\
\text{Now,finding the particular solution }\\
y_p=\frac{1}{(D'^2-2D'+4)}cosz\\
=\frac{1}{(-(1)^2-2D'+4)}cosz\\
=\frac{1}{(3-2D')}cosz\\
=\frac{3+2D'}{(9-4D'^2)}cosz\space (\text{Rationalise})\\
=\frac{3+2D'}{(9-4(-(1)^2)}cosz\\
=\frac{3+2D'}{13}cosz\\
=\frac{1}{13}(3cosz-2sinz)\\
=\frac{1}{13}(3cos(logx)-2sin(logx))\\
\text{Therefore, general solution of the given differential equation is}\\
y=x(c_1cos(\sqrt3logx)+c_2sin(\sqrt3logx))+\frac{1}{13}(3cos(logx)-2sin(logx))"